# Identifying and Improving Disability Bias in GAI-Based Resume Screening

### v1

@@ Abstract.
As Generative AI rises in adoption, its use has expanded to include domains such as hiring and recruiting. However, without examining the potential of bias, this may negatively impact marginalized populations, including people with disabilities. To address this important concern, we present a resume audit study, in which we ask ChatGPT (specifically, GPT-4) to rank a resume against the same resume enhanced with an additional leadership award, scholarship, panel presentation, and membership that are disability related. We find that GPT-4 exhibits prejudice towards these enhanced CVs. Further, we show that this prejudice can be quantifiably reduced by training a custom GPTs on principles of DEI and disability justice. Our study also includes a unique qualitative analysis of the types of direct and indirect ableism GPT-4 uses to justify its biased decisions and suggest directions for additional bias mitigation work. Additionally, since these justifications are presumably drawn from training data containing real-world biased statements made by humans, our analysis suggests additional avenues for understanding and addressing human bias.
Resume Audit, Bias, Ableism
copyright: nonejournalyear: 2024doi: conference: ; 2024; isbn: ccs: Social and professional topics People with disabilitiesccs: Social and professional topics Employment issuesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Generative Artificial Intelligence (GAI) is being increasingly used for workforce recruiting and human resource management (e.g., (Lee, 2023; Nemirovsky et al., 2021; Rane, 2023; Rathnayake and Gunawardana, 2023; Rane, 2023)). One common example is resume screening, where artificial intelligence is used to rank resumes, a task for which large language models (LLMs) such as ChatGPT are increasingly used (e.g., (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)). GAI’s advantages include optimizing the potentially time-consuming process of screening resumes to a fraction of what a purely human-driven review process would take (Tran et al., 2023), and accurately summarizing lengthy application materials to highlight a candidate’s strengths and weaknesses (Lee, 2023). However, there is a danger in using AI for hiring: AI-based resume filtering and recruitment systems are biased, for example against candidates of diverse genders (Gagandeep et al., 2023; Kodiyan, 2019) and ages (Harris, 2023). Prior work has also highlighted potential risks of disability bias in hiring (e.g., (Buyl et al., 2022; Tilmes, 2022)). However, no prior work has quantified the amount of bias due to disability in deployed AI systems.
Further, no work we are aware of has demonstrated a way to reduce disability bias in resume screening. While human involvement and collaboration has been posed as a solution to general AI-created bias (Van den Broek et al., 2021), existing research shows that even experienced recruiters with expressed skepticism for AI-based solutions may default to accepting AI-based feedback when receiving inconsistent recommendations from a system (Lacroux and Martin-Lacroux, 2022). This highlights the importance of addressing bias in the AI systems themselves, in addition to any human interventions.
This article addresses these gaps by quantifying, and then reducing, bias in AI-based resume screening specifically for people with disabilities. Disabled people, of whom there are 42.5 million in the United States (Leppert and Schaeffer, 2023), already face significant barriers to employment, including fewer callbacks and inequality in the labor market (Ameri et al., 2015; Bjørnshagen and Ugreninov, 2021). Any ableist(Cherney, 2011) bias in AI-based hiring systems could exacerbate such employment barriers. Yet these systems are already in use (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023). It is urgent that we understand, evaluate, and mitigate bias present in GAI-based resume screening for people with disabilities. Therefore, this work seeks to address the following research questions:
@@
Does a GAI-based resume screening exhibit bias against resumes that mention disability compared to those that do not? Is this bias different depending on the type of disability mentioned in the resume?
@@
Does a GAI trained on DEI principals exhibit reduced bias in comparison to a generic GAI?
@@
Do GAI explanations of rankings provide evidence for potential sources or types of bias?
After summarizing related work on bias in AI-based hiring, and quantification of bias in hiring through resume audits in Related Work (Section 2), we describe our mixed-methods resume audit study method in Methods (Section 3). We test for bias by asking ChatGPT, and a DEI and disability-justice (Berne et al., 2018) trained custom GPT, to complete a series of ranking tasks comparing a control resume to a resume enhanced with a disability-related leadership award, scholarship, panel presentation and organizational membership. We vary the type of disability mentioned in the enhanced resume, and ask both GAIs to rank each control/enhanced pair 10 times, providing an explanation each time. Our findings quantitatively demonstrate statistically significant bias in ranking, significant differences in amount of bias across different disabilities, and that training can significantly reduce bias (Section 4). Further, our qualitative analysis of GPT’s explanations for its rankings (Section 4.4) demonstrates both direct and indirect ableist reasoning.
To summarize, our work takes a systematic approach to quantify ableist bias and complement these findings with qualitative evidence. Our study method is novel because of its use of direct comparison, since our use of GPT allows ranking a standard resume against a resume enhanced with disability-related items, instead of simply measuring callbacks or responses to a single resume. Our results are especially important because we demonstrates bias using AI tools that are currently being used to rank resumes. Further, our study is the first resume audit study to uncover the reasoning behind such biases, since our use of GPT also supports collection and analysis of the rationale behind the rankings. Since “…society’s racism, misogyny, ableism, etc., tend to be overrepresented in training data …[an LLM] that has been trained on such data will pick up these kinds of problematic associations” (Bender et al., 2021). Thus, our qualitative analysis of GAI-produced text may help to uncover biased reasoning that also impacts human judgment, something rarely part of resume audit studies. Based on these findings, we highlight important avenues for further work in Discussion and Recommendations (Section 5) and Ethical Guidance (Section 7). Our recommendations could pave the path for future efforts to mitigate bias and make GAI-based recruiting systems truly useful in equitable hiring.
@@ 2. Related Work
Hiring bias is an unfortunate reality, and has been linked to the unconscious and sometimes conscious mental process which influences the evaluation of candidates (Harvard Business School, 2020), including biases based on factors such as gender, race, and ethnicity (e.g., (Kessler et al., 2019; Ock, 2022; Thanasombat and Trasviña, 2005; Zarb, 2022)). Studies have shown that aspects of a hiring profile, such as the applicant’s name, can indicate an applicant’s ethnicity and trigger a biased response (Thanasombat and Trasviña, 2005). From the perspective of a disabled job-seeker, bias is an unfortunate reality that is sometimes mitigated by controlling when, whether and how they disclose their disabilities (e.g., (Alexandrin et al., 2008; Marshall et al., 2020; Evans, 2019; Lyons et al., 2018; Grimes et al., 2019; Charmaz, 2010; Kulkarni, 2022)). Some studies propose addressing bias by raising awareness of unconscious bias and its implications, or implementing methods to classify candidates in a way that minimizes the impact of bias (Harvard Business School, 2020; Bohnet et al., 2020).
@@
The use of AI in hiring has many potential benefits, from actionable and constructive feedback for job seekers (Nemirovsky et al., 2021; Du et al., 2023) to time savings for recruiters (Tran et al., 2023). Generative AI in particular can quickly summarize and highlight important aspects of applications (Tran et al., 2023; Lee, 2023). However, AI has been demonstrated to reproduce human biases, spurring a movement in some countries such as the EU AI Act, which monitors its use for critical areas such as employment (Parliament, 2023). In the U.S., widely available AI systems are regularly used in candidate tracking with limited oversight. The emergence of low-cost GAI tools such as ChatGPT, which are in use today recruiting and hiring (e.g., (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)), has created an urgent need to understand the ethics and risks of such systems. Such risks have been documented in systems pre-dating and post-dating the advent of readily available generative AI, and we highlight some of the pertinent concerns in Section 2.1, showing that very little is known about disability bias in this context. One well understood way of quantifying bias is a resume audit study. Such studies have been traditionally used to measure human bias (e.g., biases due to racial identity (Bertrand and Mullainathan, 2004; Kang et al., 2016; Oreopoulos and Dechief, 2012), degree of ethnic identification (Derous and Ryan, 2012), queer identity and participation in LGBTQ+ organizations (Mishel, 2016), and disability status (l’Horty et al., 2022; Lippens et al., 2023)) but have also been used to measure AI bias in resume screening (Rhea et al., 2022). Section 2.2 introduces the method and summarizes some relevant findings, highlighting that although disability has been studied in a resume audit (Ameri et al., 2015), this has not yet translated into resume audits of AI (Rhea et al., 2022).
@@ 2.1. AI/ML Hiring Tools and Bias
Even before the wide-spread adoption of generative AI, AI and machine learning were widely employed in hiring, and widely studied due to concerns about bias (Lacroux and Martin-Lacroux, 2022; Pena et al., 2020; Deshpande et al., 2020; Mujtaba and Mahapatra, 2019). These biases are thought to exist because the datasets for the models carry human biases themselves (Deshpande et al., 2020). Existing research explores the different dimensions of bias present in these AI models and the tools that use them, such as biases based on socio-linguistic ability, age, gender, and race (Deshpande et al., 2020; Harris, 2023; Vaishampayan et al., 2023). It was also found that the AI/ML models could discern characteristics of a person from their resume when details weren’t explicitly given (Pena et al., 2020). This has led to efforts in reducing the bias by these algorithms through masking characteristics in resumes or creating a more human-centered AI algorithm or tool design (Vaishampayan et al., 2023; Gagandeep et al., 2023).
With the advent of generative AI and its pervasive issues with bias and ethical concerns (McDuff et al., 2019; Smith and Williams, 2021; Sheng et al., 2019, 2021), it is imperative that we revisit the question of bias and how to reduce it. Bias has been demonstrated in GAI-generated representations of a variety of minoritized identities, including people with disabilities (Zhou et al., [n. d.]; Gadiraju et al., 2023; Glazko et al., 2023). For example, LLMs have, in some contexts, associated disability with negativity (Venkit et al., 2022), or with ableist stereotypes and tropes (Gadiraju et al., 2023; Glazko et al., 2023). Few prior works have studied disability barriers in AI-based hiring systems. Two articles analyze AI-enhanced hiring processes from a technical perspective (Buyl et al., 2022; Tilmes, 2022), identifying potential concerns relating to AI, fairness and disability. Nugent et al. ((2022)) explore the concerns of disabled job seekers regarding AI in hiring and find that many parts of the hiring process, including resume screenings, can unfairly penalize disabled candidates. Kassier et al. ((2023)) study a real-world deployment of fair machine learning models which score candidates on an interactive hiring assessment, i.e. models designed to mitigate disparate impact. They compared outcomes over a data set of 400,000 people for candidates who used colorblindness, dyslexia, or ADHD accommodations to those who did not and found that the mitigation methods used were effective. However, to our knowledge, no work has empirically quantified the impact of biases against disabled jobseekers when generative AI is used in the early screening process for job seekers.
@@ 2.2. Resume Audit Studies for Uncovering Bias
A resume audit is a common method for quantifying discrimination in the hiring process (Gaddis, 2018). Such studies use deception to avoid the potential self-correction of bias by the person judging the resumes (e.g., (Kang et al., 2016; Mishel, 2016; Derous and Ryan, 2012)). More specifically, a resume audit study typically modifies an identity marker in a resume unrelated to a person’s qualifications for the job, and then measures the hireability of the job seeker, as represented by how the resume is ranked or otherwise evaluated. For example, one study modified the name at the top of a resume from “Emily” or “Greg” to “Lakisha” or “Jamal” and submitted them to real-world advertisements found in the newspaper, measuring the number of callbacks. The authors found that “white” names received 50% more callbacks (Bertrand and Mullainathan, 2004). Demonstrating bias is easiest when only one small thing (such as the name) is varied; as a result these studies typically ask a different person to look at each resume. This makes it hard to ask questions about why one resume is preferred over another.
Resume audits have been instrumental in quantifying various forms of bias, including racial discrimination influenced by names and experiences on resumes signaling different racial identities (Bertrand and Mullainathan, 2004; Kang et al., 2016; Oreopoulos and Dechief, 2012), ethnic identification levels (Derous and Ryan, 2012), LGBTQ+ identity and engagement with related organizations (Mishel, 2016), disability status (l’Horty et al., 2022; Lippens et al., 2023; Ameri et al., 2018), and even the impact of prolonged unemployment (Ghayad, 2013). Emerging work has sought to understand whether the same bias present in resume audit studies conducted on humans is present in state-of-the-art LLMs and GAI tools such as GPT. Veldanda et al. ((2023)) conducted a resume audit evaluating race, gender, political orientation, and pregnancy status, comparing the performance of Claude, Bard, and GPT when asked whether a resume modified to disclose identity was appropriate for a job category (yes or no answer). The study found limited bias across political views and pregnancy but not race and gender.

### v2

# Identifying and Improving Disability Bias in GPT-Based Resume Screening

@@ Abstract.
As Generative AI rises in adoption, its use has expanded to include domains such as hiring and recruiting. However, without examining the potential of bias, this may negatively impact marginalized populations, including people with disabilities. To address this important concern, we present a resume audit study, in which we ask ChatGPT (specifically, GPT-4) to rank a resume against the same resume enhanced with an additional leadership award, scholarship, panel presentation, and membership that are disability-related. We find that GPT-4 exhibits prejudice towards these enhanced CVs. Further, we show that this prejudice can be quantifiably reduced by training a custom GPTs on principles of DEI and disability justice. Our study also includes a unique qualitative analysis of the types of direct and indirect ableism GPT-4 uses to justify its biased decisions and suggest directions for additional bias mitigation work. Additionally, since these justifications are presumably drawn from training data containing real-world biased statements made by humans, our analysis suggests additional avenues for understanding and addressing human bias.
Resume Audit, Bias, Ableism, GPT
journalyear: 2024copyright: rightsretainedconference: The 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: The 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3658933isbn: 979-8-4007-0450-5/24/06ccs: Social and professional topics People with disabilitiesccs: Social and professional topics Employment issuesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Generative Artificial Intelligence (GAI) is being increasingly used for workforce recruiting and human resource management (e.g., (Lee, 2023; Nemirovsky et al., 2021; Rane, 2023; Rathnayake and Gunawardana, 2023)). One common example is resume screening, where artificial intelligence is used to rank resumes, a task for which the use of large language models (LLMs) such as ChatGPT is becoming more frequently discussed (e.g., (Lever Team, 2023; Jennings, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)). GAI’s advantages include optimizing the potentially time-consuming process of screening resumes to a fraction of what a purely human-driven review process would take (Tran et al., 2023), and accurately summarizing lengthy application materials to highlight a candidate’s strengths and weaknesses (Lee, 2023). However, there is a danger in using AI for hiring: AI-based resume filtering and recruitment systems are biased, for example against candidates of diverse genders (Gagandeep et al., 2023; Kodiyan, 2019) and ages (Harris, 2023). Prior work has also highlighted potential risks of disability bias in hiring (e.g., (Buyl et al., 2022; Tilmes, 2022)). Yet, no prior work has quantified the amount of bias due to disability when using popular GAI tools such as ChatGPT for resume screenings and candidate summaries.
Further, no work we are aware of has demonstrated a way to reduce disability bias in GAI resume screening. While human involvement and collaboration has been posed as a solution to general AI-created bias (Van den Broek et al., 2021), existing research shows that even experienced recruiters with expressed skepticism for AI-based solutions may default to accepting AI-based feedback when receiving inconsistent recommendations from a system (Lacroux and Martin-Lacroux, 2022). This highlights the importance of addressing bias in the AI systems themselves, in addition to any human interventions.
This article addresses these gaps by quantifying, and then reducing, bias in GAI-based resume screening specifically for people with disabilities. Disabled people, of whom there are 42.5 million in the United States (Leppert and Schaeffer, 2023), already face significant barriers to employment, including fewer callbacks and inequality in the labor market (Ameri et al., 2018; Bjørnshagen and Ugreninov, 2021). Any ableist (Cherney, 2011) bias in AI-based hiring systems could exacerbate such employment barriers. Yet these systems are already in use (Lever Team, 2023; Jennings, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023). It is urgent that we understand, evaluate, and mitigate bias present in GAI-based resume screening for people with disabilities. Therefore, this work seeks to address the following research questions:
@@
Does a GPT-based resume screening exhibit bias against resumes that mention disability compared to those that do not? Is this bias different depending on the type of disability mentioned in the resume?
@@
Does a GPT trained on DEI principals exhibit reduced bias in comparison to a generic GPT?
@@
Do GPT explanations of rankings provide evidence for potential sources or types of bias?
After summarizing related work on bias in AI-based hiring and quantification of bias in hiring through resume audits in Related Work (Section 2), we describe our mixed-methods resume audit study method in Methods (Section 3). We test for bias by asking ChatGPT, and a DEI and disability-justice (Berne et al., 2018) trained custom GPT, to complete a series of ranking tasks comparing a control resume to a resume enhanced with a disability-related leadership award, scholarship, panel presentation, and organizational membership. We vary the type of disability mentioned in the enhanced resume, and ask both GAIs to rank each control/enhanced pair ten times, providing an explanation each time. Our findings quantitatively demonstrate bias in ranking, differences in the amount of bias across different disabilities, and that training can reduce bias (Section 4). Further, our qualitative analysis of GPT’s explanations for its rankings (Section 4.4) demonstrates both direct and indirect ableist reasoning.
To summarize, our work takes a systematic approach to quantify ableist bias and complement these findings with qualitative evidence. Our study method is novel because of its use of direct comparison, since our use of GPT allows ranking a standard resume against a resume enhanced with disability-related items, instead of simply measuring callbacks or responses to a single resume. Our results are especially important because we demonstrate bias using popular GAI tools that are currently being used to rank resumes. Further, our study is the first resume audit study to uncover the reasoning behind such biases, since our use of GPT also supports the collection and analysis of the rationale behind the rankings. Since “…society’s racism, misogyny, ableism, etc., tend to be overrepresented in training data …[an LLM] that has been trained on such data will pick up these kinds of problematic associations” (Bender et al., 2021). Thus, our qualitative analysis of GAI-produced text may help to uncover biased reasoning that also impacts human judgment, something that is rarely part of resume audit studies. Based on these findings, we highlight important avenues for further work in Discussion and Recommendations (Section 5) and Ethical Guidance (Section 7). Our recommendations could pave the path for future efforts to mitigate bias and make GAI-based recruiting systems truly useful in equitable hiring.
@@ 2. Related Work
Hiring bias is an unfortunate reality, and has been linked to the unconscious and sometimes conscious mental process that influences the evaluation of candidates (Harvard Business School, 2020), including biases based on factors such as gender, race, and ethnicity (e.g., (Kessler et al., 2019; Ock, 2022; Thanasombat and Trasviña, 2005; Zarb, 2022)). Studies have shown that aspects of a hiring profile, such as the applicant’s name, can indicate an applicant’s ethnicity and trigger a biased response (Thanasombat and Trasviña, 2005). From the perspective of a disabled job-seeker, bias is an unfortunate reality that is sometimes mitigated by controlling when, whether, and how they disclose their disabilities (e.g., (Alexandrin et al., 2008; Marshall et al., 2020; Evans, 2019; Lyons et al., 2018; Grimes et al., 2019; Charmaz, 2010; Kulkarni, 2022)). Some studies propose addressing bias by raising awareness of unconscious bias and its implications, or implementing methods to classify candidates in a way that minimizes the impact of bias (Harvard Business School, 2020; Bohnet et al., 2020).
@@
The use of AI in hiring has many potential benefits, from actionable and constructive feedback for job seekers (Nemirovsky et al., 2021; Du et al., 2023) to time savings for recruiters (Tran et al., 2023). Generative AI, in particular, can quickly summarize and highlight important aspects of applications (Tran et al., 2023; Lee, 2023). However, AI has been demonstrated to reproduce human biases, spurring a movement in some countries such as the EU AI Act, which monitors its use for critical areas such as employment (Parliament, 2023). In the U.S., widely available AI systems are regularly used in candidate tracking with limited oversight. The emergence of low-cost GAI tools such as ChatGPT, which are in use today in recruiting and hiring (e.g., (Lever Team, 2023; Jennings, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)), has created an urgent need to understand the ethics and risks of such systems. Such risks have been documented in systems pre-dating and post-dating the advent of readily available GAI, and we highlight some of the pertinent concerns in Section 2.1, showing that very little is known about disability bias in this context. One well-understood way of quantifying bias is a resume audit study. Such studies have been traditionally used to measure human bias (e.g., biases due to racial identity (Bertrand and Mullainathan, 2004; Kang et al., 2016; Oreopoulos and Dechief, 2012), degree of ethnic identification (Derous and Ryan, 2012), queer identity and participation in LGBTQ+ organizations (Mishel, 2016), and disability status (l’Horty et al., 2022; Lippens et al., 2023)) but have also been used to measure AI bias in resume screening (Rhea et al., 2022). Section 2.2 introduces the method and summarizes some relevant findings, highlighting that although disability has been studied in a resume audit (Ameri et al., 2018), this has not yet translated into resume audits of AI (Rhea et al., 2022; Veldanda et al., 2023; Walker, 2024).
@@ 2.1. AI/ML Hiring Tools and Bias
Even before the widespread adoption of generative AI, AI and machine learning were widely employed in hiring, and widely studied due to concerns about bias (Lacroux and Martin-Lacroux, 2022; Pena et al., 2020; Deshpande et al., 2020; Mujtaba and Mahapatra, 2019). These biases are thought to exist because the datasets for the models carry human biases themselves (Deshpande et al., 2020). Existing research explores the different dimensions of bias present in these AI models and the tools that use them, such as biases based on socio-linguistic ability, age, gender, and race (Deshpande et al., 2020; Harris, 2023; Vaishampayan et al., 2023; Mei et al., 2023). It was also found that the AI/ML models could discern characteristics of a person from their resume when details weren’t explicitly given (Pena et al., 2020). This has led to efforts to reduce the bias in these algorithms through masking characteristics in resumes or creating a more human-centered AI algorithm or tool design (Vaishampayan et al., 2023; Gagandeep et al., 2023).
With the advent of generative AI and its pervasive issues with bias and ethical concerns (McDuff et al., 2019; Smith and Williams, 2021; Sheng et al., 2019, 2021; Mei et al., 2023), it is imperative that we revisit the question of bias and how to reduce it. Bias has been demonstrated in GAI-generated representations of a variety of minoritized identities (Mei et al., 2023), including people with disabilities (Zhou et al., 2023; Gadiraju et al., 2023; Glazko et al., 2023). For example, LLMs have, in some contexts, associated disability with negativity (Venkit et al., 2022) or with ableist stereotypes and tropes (Gadiraju et al., 2023; Glazko et al., 2023). Few prior works have studied disability barriers in AI-based hiring systems. Two articles analyze AI-enhanced hiring processes from a technical perspective (Buyl et al., 2022; Tilmes, 2022), identifying potential concerns relating to AI, fairness, and disability. Nugent et al. ((2022)) explore the concerns of disabled job seekers regarding AI in hiring and find that many parts of the hiring process, including resume screenings, can unfairly penalize disabled candidates. Kassier et al. ((2023)) study a real-world deployment of fair machine learning models that score candidates on an interactive hiring assessment, i.e. models designed to mitigate disparate impact. They compared outcomes over a data set of 400,000 people for candidates who used colorblindness, dyslexia, or ADHD accommodations to those who did not and found that the mitigation methods used were effective. However, to our knowledge, no work has empirically quantified the impact of biases against disabled jobseekers when generative AI is used in the early screening process for job seekers.
@@ 2.2. Resume Audit Studies for Uncovering Bias
A resume audit is a common method for quantifying discrimination in the hiring process (Gaddis, 2018). Such studies use deception to avoid the potential self-correction of bias by the person judging the resumes (e.g., (Kang et al., 2016; Mishel, 2016; Derous and Ryan, 2012)). More specifically, a resume audit study typically modifies an identity marker in a resume unrelated to a person’s qualifications for the job, and then measures the hireability of the job seeker, as represented by how the resume is ranked or otherwise evaluated. For example, one study modified the name at the top of a resume from “Emily” or “Greg” to “Lakisha” or “Jamal” and submitted them to real-world advertisements found in the newspaper, measuring the number of callbacks. The authors found that “white” names received 50% more callbacks (Bertrand and Mullainathan, 2004). Demonstrating bias is easiest when only one small thing (such as the name) is varied; as a result, these studies typically ask a different person to look at each resume. This makes it hard to ask questions about why one resume is preferred over another.
Resume audits have been instrumental in quantifying various forms of bias, including racial discrimination influenced by names and experiences on resumes signaling different racial identities (Bertrand and Mullainathan, 2004; Kang et al., 2016; Oreopoulos and Dechief, 2012), ethnic identification levels (Derous and Ryan, 2012), LGBTQ+ identity, and engagement with related organizations (Mishel, 2016), disability status (l’Horty et al., 2022; Lippens et al., 2023; Ameri et al., 2018), and even the impact of prolonged unemployment (Ghayad, 2013). Emerging work has sought to understand whether the same bias present in resume audit studies conducted on humans is present in state-of-the-art LLMs and GAI tools such as GPT (Walker, 2024; Veldanda et al., 2023). Veldanda et al. ((2023)) conducted a resume audit evaluating race, gender, political orientation, and pregnancy status, comparing the performance of Claude, Bard, and GPT when asked whether a resume modified to disclose identity was appropriate for a job category (yes or no answer). The study found limited bias across political views and pregnancy but not race and gender (Veldanda et al., 2023).